Please explain how to solve this?
-5x-10=2-(x+4)
thank-you

Answer by oscargut(2103)

-5x-10=2-(x+4)
then
-5x-10=2-x-4
then
-5x+x=2-4+10
-4x=8
x=8/(-4)=-2

Solution: x=-2